OpenStudy (anonymous):

3(5z-7)-2(9z-11)=4(8z-13)-17

OpenStudy (shadowfiend):

Are you trying to solve for z?

OpenStudy (shadowfiend):

Hm. I would assume as much, but it's not exactly clear.

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

15z-21-18z+22 = 32z-52-17 15z-18z+22-21 = 32z-69 -3z+1 = 32z-69 1+69 = 32z+3z 70 = 35z 35z = 70 z = 70/35 z= 2